I need to know the real and imaginary parts of the complex number.
(17+2i)i^47. All I know is that I need to work the problem out before knowing what each part is.

Answer by venugopalramana(3286)

I need to know the real and imaginary parts of the complex number.
(17+2i)i^47.
=(17+2i){(i^2)^23}*i=(17+2i)(-1)^23*i
=-(17+2i)i
=-17i-2
R.P=-2
I.P.=-17
All I know is that I need to work the problem out before knowing what each part is.